I'm very confused! Can you help me with this?
Thank you!
two solutions for you to create a quadratic equation:
1. x=1 x=5
2. x=5 x=0

Answer by checkley77(12844)

1. x=1 x=5
(X^2-6X+5)=0
(X-5)(X-1)=0
X-5=0
X=5
X-1=0
X=1
-----------------------
2. x=5 x=0
X^2-5X=0
X(X-5)=0
X=0
X-5=0
X=5